# JigsawComm: Joint Semantic Feature Encoding and Transmission for Communication-Efficient Cooperative Perception

### v1

@@ 3.3 Feature Utility Estimator (FUE)
Utility Proxy. To address C1, we propose to maximize a utility function $U(\cdot)$. It represents the total contribution of all agents’ transmitted data, expressed as the sum of individual agent’s feature utility minus their pairwise redundancy. Essentially, our utility function can be regarded as a proxy for the overall information about the CP inference result (on the fused feature) that is contained within the transmitted data. The greater this mutual information is, the higher the perception accuracy [qin2021semantic, strinati2021_6g].
@@
Here, $m_{i}^{l}\in\{0,1\}$ is the binary transmission decision for $\mathbf{f}_{i}^{l}$, $u_{i}^{l}$ is its estimated importance, representing the amount of information contained in $i$’s feature at location $l$ about the CP inference result. Also, $\tilde{u}_{ij}^{l}$ approximates the pairwise redundancy, upper-bounding the shared information that $i$ and $j$’s features both contain about the CP inference result, where the $\frac{1}{2}$ factor is to avoid double-counting. Finally, the frame-level utility is $U=\sum_{l=1}^{L}U^{l}$.
FUE Head. To estimate the feature utility, we implement FUE as a pointwise linear decoder, similar to the architecture of a classification or regression head $u_{i}^{l}\;=\;\mathrm{ReLU}\!\big(\mathbf{w}^{\top}\mathbf{f}_{i}^{l}+b\big)$, resulting in a sparse meta utility map $\mathbf{U}_{i}\in\mathbb{R}^{1\times H\times W}$. The layer is vectorized as a $1{\times}1$ convolution for efficiency and interpretability, since each location already contains neighboring cells’ information from the convolutional layers in the encoder. The FUE is end-to-end trained with the transmission policy described in Sec. 3.4 and aligned with the perception accuracy loss. Therefore, its value directly relates to the transmission decision such that higher utility values correlate with higher transmission priority and larger contribution to the accuracy.
@@ 3.4 Transmission Scheduler
Even after transforming the optimization into maximizing the learned utility proxy, the challenges of non-differentiability in scheduling (C2) and the NP-hard combinatorial knapsack problem (C3) still exist. To solve these, we design a scheduling algorithm that is (i) differentiable at training and (ii) deterministic at inference with provably optimal guarantee given fixed transmission cost per feature.
@@
This policy can be deterministically computed by all agents, making it practically compatible with both centralized (V2I) and decentralized (V2V) operation (Sec. 5).
@@
Train-Time Differentiable Scheduling. To address C2, we relax the discrete selection during training. Instead, we model the mask $m_{i}^{l}$ as the composition of two gates:
@@
(2) Cross-agent top-1 (categorical, per cell). To model the “pick one agent per cell” constraint differentiably, we apply the Gumbel–Softmax reparameterization [jang2016categorical] on the vector of scores $\mathbf{u}^{l}=(u_{1}^{l},\dots,u_{N}^{l})$ with an independent Gumbel noise per agent: $g_{i}^{l}\sim\mathrm{Gumbel}(0,1)$ with $g_{i}^{l}=-\log(-\log\varepsilon_{i}^{l})$, $\varepsilon_{i}^{l}\sim\mathrm{Uniform}(0,1)$ , and a temperature $\gamma>0$, resulting in:
@@
In the backward pass, we compute gradients as if the mask was the soft, differentiable approximation $\tilde{m}_{i}^{l}=\alpha_{i}^{l}\boldsymbol{\beta}_{i}^{l}(\gamma)$. $\gamma$ follows the same annealing schedule as $\eta$ during training.
@@
Proposition 1 (Consistency of differentiable and deterministic schedulers).
@@
Prop. (1) indicates that the train-time relaxed mask converges pointwise to the deterministic policy. Also, instead of a heuristic design, we show that this policy is provably optimal for the utility proxy in Eq. (4) and the knapsack problem under uniform costs.
Theorem 1 (Singleton optimality).
Proof.
Figure 4: Visualization of feature map examples.
Theorem (1) ensures that there is at most one agent with the highest utility to transmit per location cell, inherently eliminating cross-agent redundancy, as visualized in Fig. 3. This reduces the knapsack problem (C3) of selecting features across both agent and location dimensions to selecting the best set of location cells to transmit. Although the problem remains a knapsack problem, it becomes a uniform-cost knapsack, since the number of channels is fixed per feature, resulting in the same transmission cost. It is known that the uniform-cost knapsack reduces to a cardinality-constrained selection, which can be optimally solved by taking the elements with the largest utilities [kellerer2004multidimensional], as shown in Theorem (2).
Theorem 2 (Greedy optimality under equal costs).
Proof.
By Theorem (1), each cell yields at most one admissible candidate; thus the global problem is
$\max_{\mathcal{S}\subseteq\mathcal{C}}\ \sum_{(i,l)\in\mathcal{S}}u_{i}^{l}\quad\text{s.t. }\ |\mathcal{S}|\leq K,$ (11)
where $\mathcal{C}$ is the set of at-most-one-per-cell candidates and all costs equal $c$. This is a cardinality-constrained selection problem. Let $\mathcal{G}$ be the greedy solution: the $K$ items with the largest utilities. Assume for contradiction that $\mathcal{G}$ is not optimal. Let $\mathcal{O}$ be an optimal solution with $|\mathcal{O}|\leq K$ and $\sum_{o\in\mathcal{O}}u(o)>\sum_{g\in\mathcal{G}}u(g)$. If $|\mathcal{O}|<K$, augment $\mathcal{O}$ by adding the highest-utility items not in $\mathcal{O}$ until its size reaches $K$; this cannot decrease its objective value because utilities are nonnegative, so w.l.o.g. take $|\mathcal{O}|=K$.
Order items in $\mathcal{G}$ and $\mathcal{O}$ so that $u(g_{1})\geq\cdots\geq u(g_{K})$ and $u(o_{1})\geq\cdots\geq u(o_{K})$. Since $\mathcal{G}$ contains the top-$K$ utilities overall, we must have $u(g_{j})\geq u(o_{j})$ for every $j=1,\dots,K$. Summing over $j$ yields $\sum_{j=1}^{K}u(g_{j})\geq\sum_{j=1}^{K}u(o_{j})$, contradicting $\sum_{o\in\mathcal{O}}u(o)>\sum_{g\in\mathcal{G}}u(g)$. Hence $\mathcal{G}$ is optimal. ∎
@@ 3.5 End-to-End Communication-Aware Training
We unify these components into an end-to-end trainable framework. The final loss combines the downstream task loss (e.g. focal, cross-entropy, etc.) with the semantic loss:
@@
where $\lambda$ is a hyperparameter balancing the sparsity of the encoded feature and task accuracy. The $\mathcal{L}_{\text{task}}$ gradient flows through the differentiable scheduler (Sec. 3.4) to train the FUE head (Sec. 3.3), while $\mathcal{L}_{\text{semantic}}$ shapes the encoder (Sec. 3.2). By learning the sparsity threshold $\kappa$ and the utility threshold $\tau$ jointly, JigsawComm learns to activate and transmit features only when they contribute to CP accuracy, maximizing the utility of every transmitted bit, as shown in Fig. 4.
@@ 4 Evaluation
Dataset OPV2V DAIR-V2X
Model mAP @0.5 mAP @0.7 Per-Agent Size (KB) Total Data Size (KB) Rel. Eff. (%) mAP @0.5 mAP @0.7 Per-Agent Size (KB) Total Data Size (KB) Rel. Eff. (%)
No Fusion 0.68 0.60 - - - 0.50 0.44 - - -
F-Cooper 0.89 0.79 8,800.00 22,792.00 0.19 0.56 0.39 16,384.00 32,768.00 0.70
V2X-ViT 0.87 0.67 8,800.00 22,792.00 0.18 0.68 0.50 16,384.00 32,768.00 0.84
AttFusion()(*) 0.91 0.82 8,800.00 22,792.00 0.19 0.65 0.51 16,384.00 32,768.00 0.81
()(*) + Enc-Dec 0.89 0.81 4,400.00 11396.00 0.38 0.65 0.51 8,192.00 16,384.00 1.61
Where2comm 0.85 0.60 1,152.47 2,794.72 1.47 0.46 0.23 14,692.88 29,385.76 0.64
JigsawComm 0.92 0.82 17.46 43.84 100.00 0.69 0.52 148.96 281.97 100.00
@@ 4.1 Experimental Setup
Dataset. We evaluate JigsawComm on two CP datasets: (i) OPV2V [xu2022opv2v], which is a large-scale benchmark collected across 70 diverse scenes from 8 towns in the digital twin simulator CARLA [Dosovitskiy17Carla]. It includes 11,464 frames of LiDAR point cloud data from 2-5 agents and 232,913 annotated 3D bounding boxes. (ii) DAIR-V2X [yu2022dair], which is the first large-scale real-world CP dataset for vehicle-infrastructure cooperation, consisted of 71,254 frames.
Implementation Details. We adopt PointPillars [lang2019pointpillars] as the LiDAR BEV backbone and a max-out module as the fusion backbone [f_cooper]. We train JigsawComm for up to 100 epochs using Adam ($\beta_{1}\!=\!0.9,\beta_{2}\!=\!0.999$), with a base learning rate of 0.002. The Gumbel-Softmax temperatures $\eta,\gamma$ are annealed from $0.9\to 0.1$. We select $\lambda=1000$ for OPV2V and $\lambda=100$ for DAIR-V2X, through standard cross validation.
@@
Baselines. We follow the evaluation protocol that has become standard in communication-efficient CP [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x]: (i) a no-fusion single-vehicle detector, (ii) strong full-transmission intermediate-fusion methods that transmit dense BEV features, including F-Cooper [f_cooper], V2X-ViT [xu2022v2x], and AttFusion [xu2022opv2v], and (iii) communication-efficient CP via Where2comm [hu2022where2comm], which is the de facto canonical communication-efficient baseline adopted by recent works [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x]. Although some recent approaches [hu2024communication, mao2025diffcp, zhao2025quantv2x] achieve comparable bandwidth requirements, they suffer from practical limitations in generalizability or computation latency, as discussed in Sec. 2.2.
@@ 4.2 Quantitative Results
[…]
@@
Figure 5: Tradeoff between accuracy and required bandwidth.
Accuracy-Bandwidth Tradeoff. Figure 5 visualizes JigsawComm’s performance versus the required communication bandwidth ($BW$). The baselines are fixed points, representing inflexible, inefficient strategies. JigsawComm’s curve is generated by varying the communication budget $B$. The accuracy of JigsawComm saturates quickly after $BW\geq$1,000 Kbps, making it a flexible, generalizable, and pragmatic solution for deployment under the limited bandwidth offered by V2X (e.g, $\leq$20 Mbps) [FCC_DSRC, CV2X_ETSI_EN_303_613].
Figure 6: Accuracy, feature size, and required bandwidth grouped by agent counts on OPV2V. Instead of a linear increase in total feature size with more agents, JigsawComm induces an asymptotic $\mathcal{O}(1)$ communication cost while achieving high accuracy.
Scalability with Number of Agents. JigsawComm achieves an asymptotic $\mathcal{O}(1)$ communication cost, where the total bandwidth is bounded by the scene’s information content in a single sparse feature, not the number of participating agents $N$. We validate this claim empirically in Fig. 6 using the OPV2V dataset, which includes 2 to 5 agents. JigsawComm’s accuracy remains high as the number of agents varies. The total feature size grows slightly from $\approx$40 KB at $N=2$ to $\approx$58 KB at $N=4$, which is attributed to the fact that more agents increases collective scene coverage. Our FUE correctly identifies these non-redundant cells as having high utility and adds them to the ‘jigsaw puzzle’.
Since OPV2V is limited to 5 agents, we conduct further experiments on a custom-simulated dataset to analyze the asymptotic trend. We utilized the CARLA digital twin simulator [Dosovitskiy17Carla] in conjunction with the OpenCDA framework [opencda] for coordinated multi-agent simulation and data recording. We set up the scene within CARLA’s ‘Town10HD’ map with 10–20 CAVs following realistic traffic rules, at a large multi-lane intersection. The data was collected following the same format and conventions as the public OPV2V benchmark dataset, including sensor configurations, coordinate systems, and CAV parameters, to ensure compatibility with existing CP models and evaluation pipelines. Fig. 7 presents the result of JigsawComm on the custom dataset. We can observe that JigsawComm’s total communication cost does not double when scaling from 10 to 20 agents. It saturates and remains relatively flat when the number of agents is greater than 15, empirically verifying the asymptotic $\mathcal{O}(1)$ communication cost claim. As more agents join, the collective coverage saturates, the overlap becomes dominant, and the marginal contribution of each new agent becomes smaller.
@@
End-to-End Efficiency. Beyond theoretical bandwidth savings, we validate JigsawComm’s end-to-end efficiency. The average end-to-end latency is $45.82\pm 26.96$ ms, which is well within the stringent 100 ms real-time budget for autonomous driving perception [ADConstraints]. This total latency consists of $28.27\pm 23.88$ ms for computation on an RTX 4090 GPU and $17.55\pm 7.14$ ms for communication (assuming a 20 Mbps bandwidth constraint), including the total feature payload from all agents and the meta utility maps with a small total overhead of $0.08\pm 0.04$KB.

### v2

@@ 3.3 Feature Utility Estimator (FUE)
Utility Proxy. As established in task-oriented and semantic communication [qin2021semantic, strinati2021_6g], higher CP accuracy correlates with transmitted data retaining more information about the final inference result. Therefore, to address (C1), we design a utility function $U(\cdot)$ that serves as a proxy for the downstream CP accuracy and can be computed locally from the agents’ features. $U$ aggregates the individual feature contributions of all agents and penalizes pairwise redundancy. The utility at each spatial cell $l$ is defined as:
@@
Here, $m_{i}^{l}\in\{0,1\}$ is the binary transmission decision for $\mathbf{f}_{i}^{l}$ and $u_{i}^{l}$ is its estimated importance—the FUE predicted contribution of agent $i$’s feature at cell $l$ to the CP result. The term $\tilde{u}_{ij}^{l}$ conservatively upper-bounds the pairwise redundancy between agents $i$ and $j$, with the $\frac{1}{2}$ factor avoiding double-counting [cover1999elements]. Since features in the same BEV cell describe the same physical location; overlapping observations typically differ only in strength, making $\min(\cdot)$ a tractable upper bound on shared information. The frame-level utility is $U=\sum_{l=1}^{L}U^{l}$, which serves as a first-order approximation of task-relevant information by rewarding informative features and penalizing redundancy [cover1999elements].
Scope of the Utility Proxy. We emphasize that $U$ is not an exact information-theoretic quantity but a learned, task-aligned intermediate representation acting as an inductive bias that facilitates transmission scheduling: the FUE predicted utilities $u_{i}^{l}$ are trained end-to-end to correlate with CP accuracy in a data-driven fashion rather than derived from a closed-form mutual information expression, which is computationally infeasible [cover1999elements]. The theoretical guarantees in Sec. 3.4 (Theorems 3.13.2) hold with respect to this learned proxy; their practical value depends on how well $U$ tracks the true task performance, which we validate empirically in Sec. 4.
FUE Head. We implement the FUE as a single $1{\times}1$ convolution (pointwise linear decoder): $u_{i}^{l}=\mathrm{ReLU}\!\big(\mathbf{w}^{\top}\mathbf{f}_{i}^{l}+b\big)$, producing a sparse meta utility map $\mathbf{U}_{i}\in\mathbb{R}^{1\times H\times W}$. Each cell’s receptive field already spans its spatial neighborhood through the encoder’s convolutional layers, so a pointwise head suffices. The FUE is trained end-to-end through the differentiable scheduler (Sec. 3.4): the task loss gradient back-propagates through the scheduling decisions to the FUE weights, aligning higher utility values with higher transmission priority and greater contribution to accuracy.
@@ 3.4 Transmission Scheduler
Even after transforming the optimization into maximizing the learned utility proxy, the challenges of non-differentiability in scheduling (C2) and the NP-hard combinatorial knapsack problem (C3) still exist. To solve these, we design a scheduling algorithm that is $(i)$ differentiable at training and $(ii)$ deterministic at inference, with a theoretically grounded optimality guarantee under the learned utility proxy for uniform-cost features.
@@
This policy can be deterministically computed by all agents, making it practically compatible with both centralized (V2I) and decentralized (V2V) operation (Sec. 5). We next show that this policy is optimal for the learned utility proxy–i.e. it maximizes $U$ among all feasible schedules, as formalized in Theorem 3.1 and 3.2.
@@
Theorem 3.1(Singleton optimality)
Proof
Theorem 3.1 guarantees at most one agent per cell, inherently eliminating cross-agent redundancy (visualized in Fig. 3). This reduces the joint agent-and-cell selection problem (C3) to selecting a subset of cells, each with a single pre-determined agent. Because every cell carries a $C$-channel feature vector plus an index, the per-cell cost is uniform for non-zero features [hu2022where2comm], and the resulting uniform-cost knapsack is optimally solved by selecting the cells with the largest utilities [kellerer2004multidimensional], as formalized in Theorem 3.2.
Theorem 3.2(Greedy optimality under equal costs)
We provide the proof to Theorem 3.2 in Appendix Sec. 0.A. We illustrate the incorporation of the transmission policy into the end-to-end training pipeline below, which is essential in aligning the FUE-learned utility representation with the CP accuracy objective.
Train-Time Differentiable Scheduling. Addressing the non-differentiable scheduling challenge (C2), we relax the discrete selection during training. Instead, we model the mask $m_{i}^{l}$ as the composition of two gates:
@@
(2) Cross-agent top-1 (categorical, per cell). To model the “pick one agent per cell” constraint differentiably, we apply the Gumbel–Softmax reparameterization [jang2016categorical] to the score vector $\mathbf{u}^{l}=(u_{1}^{l},\dots,u_{N}^{l})$ with independent Gumbel noise $g_{i}^{l}\sim\mathrm{Gumbel}(0,1)$ and temperature $\gamma>0$:
@@
In the backward pass, we compute gradients as if the mask was the soft, differentiable approximation $\tilde{m}_{i}^{l}=\alpha_{i}^{l}\boldsymbol{\beta}_{i}^{l}(\gamma)$. $\gamma$ follows the same annealing schedule as $\eta$.
@@
Proposition 1(Consistency of differentiable and deterministic schedulers)
@@
Prop. 1 indicates that the train-time policy converges pointwise to the inference policy.
Figure 4: Feature map visualization. The semantic loss and learnable threshold $\kappa$ produce sparse features that concentrate activations on semantically relevant regions.
@@ 3.5 End-to-End Communication-Aware Training
We unify these components into a single end-to-end trainable framework. The final loss combines the downstream task loss (e.g. focal loss) with the semantic loss:
@@
where $\lambda$ balances feature sparsity and task accuracy. The $\mathcal{L}_{\text{task}}$ gradient flows through the differentiable scheduler (Sec. 3.4) to train the FUE head (Sec. 3.3), while $\mathcal{L}_{\text{semantic}}$ shapes the encoder (Sec. 3.2). By jointly learning the sparsity threshold $\kappa$ and utility threshold $\tau$, JigsawComm activates and transmits features only when they contribute to CP accuracy, as illustrated in Fig. 4. We note that a naive alternative—penalizing the transmission masks $\{\mathbf{M}_{i}\}$ directly—performs worse in practice because it penalizes all communication indiscriminately, disregarding semantic importance and redundancy.
3.6 Robustness Enhancement
Robustness to Misalignment. The top-1-per-cell policy in Eq. (7) compares utilities $u_{i}^{l}$ across agents at each spatial cell $l$. In practice, localization noise and communication delays introduce translational offsets between agents’ coordinate frames [zhang2023robust], causing the scheduler to compare misaligned cells. To address this, we introduce a lightweight alignment step that estimates and corrects offsets using only the already-exchanged meta utility maps $\{\mathbf{U}_{i}\}$, requiring no additional communication. Inspired by the feature matching and warping in [Zhang_2024_CVPR], for each non-ego agent $j$, we extract the top-$K$ cells by utility value from both $\mathbf{U}_{\text{ego}}$ and $\mathbf{U}_{j}$ as sparse keypoint sets and estimate a translational offset $\boldsymbol{\delta}_{j}$ via RANSAC [derpanis2010ransac]: in each iteration, we sample one point from each set, compute the candidate offset, and count inliers within $\varepsilon$ cells. The best hypothesis is refined by averaging inlier correspondences. Offsets with low confidence are discarded to avoid interpolation artifacts. Given the estimated offsets, each non-ego utility map $\mathbf{U}_{j}$ and features $\mathbf{F}_{j}$ are warped by $\boldsymbol{\delta}_{j}$ via bilinear grid sampling before scheduling and fusion.
Robustness to Packet Loss. The top-1-per-cell policy also provides a natural fallback under packet loss: if a non-ego agent’s scheduled feature is dropped during transmission, the ego vehicle uses its own local feature at the affected cell, requiring no additional communication. Because the ego’s feature is always available and semantically meaningful in visible areas, this fallback degrades gracefully (Appendix 0.E) and guarantees the performance lower-bounded by local detection. Alternatively, a Top-$K$ policy could pre-select backup agents when more bandwidth is available, where the scheduler pre-selects $K$ candidate agents per cell and falls back to the next-best candidate upon packet loss.
@@ 4 Evaluation
Table 1: Main results on OPV2V and DAIR-V2X. All methods transmit their full output (unlimited budget $B$). Rel. Eff. is the mAP@0.5-to-total-size ratio, normalized to JigsawComm ($100\%$). JigsawComm achieves SOTA accuracy while reducing data volume by over $500{\times}$.
OPV2V DAIR-V2X
Method mAP@.5 mAP@.7 Per-Ag. Total Rel. Eff. mAP@.5 mAP@.7 Per-Ag. Total Rel. Eff.
(KB) (KB) (%) (KB) (KB) (%)
No Fusion 0.68 0.60 0.50 0.44
F-Cooper [f_cooper] 0.89 0.79 8,800 22,792 0.19 0.56 0.39 16,384 32,768 0.70
V2X-ViT [xu2022v2x] 0.87 0.67 8,800 22,792 0.18 0.68 0.50 16,384 32,768 0.84
@@
$\ast$ + Enc-Dec 0.89 0.81 4,400 11,396 0.38 0.65 0.51 8,192 16,384 1.61
Where2comm [hu2022where2comm] 0.85 0.60 1,152 2,795 1.47 0.59 0.46 6,295 12,590 1.91
ERMVP [Zhang_2024_CVPR] 0.89 0.79 845 2,189 1.97 0.64 0.51 2,914 5,827 4.47
CoST [tang2025cost] 0.87 0.76 4,194 12,163 0.35 0.62 0.47 4,194 8,389 3.01
[…]
@@ 4.1 Experimental Setup
Datasets. We evaluate JigsawComm on two CP benchmarks: (i) OPV2V [xu2022opv2v], a large-scale dataset with 11,464 LiDAR frames from 2–5 agents across 70 scenes in 8 CARLA [Dosovitskiy17Carla] towns; and (ii) DAIR-V2X [yu2022dair], the first large-scale real-world CP dataset for vehicle-infrastructure cooperation, containing 71,254 frames.
Implementation Details. We adopt PointPillars [lang2019pointpillars] as the LiDAR BEV backbone and a max-out module as the fusion backbone [f_cooper]. We train JigsawComm for up to 100 epochs using Adam ($\beta_{1}\!=\!0.9,\beta_{2}\!=\!0.999$), with a base learning rate of 0.002. The Gumbel-Softmax temperatures $\eta,\gamma$ are annealed from $0.9\to 0.1$.
@@
Baselines. We follow the standard evaluation protocol in communication-efficient CP [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x] and evaluate: (i) a no-fusion single-vehicle detector; (ii) full-transmission intermediate-fusion (F-Cooper [f_cooper], V2X-ViT [xu2022v2x], AttFusion [xu2022opv2v]); and (iii) communication-efficient methods, including the canonical baseline Where2comm [hu2022where2comm] and prior art ERMVP (CVPR’24 [Zhang_2024_CVPR]) and CoST (ICCV’25 [tang2025cost]). All baselines are reproduced using their official codebases with the PointPillars backbone [lang2019pointpillars] for a fair comparison.
@@ 4.2 Quantitative Results
Main Results. Table 1 presents our main comparison. On OPV2V, JigsawComm matches or exceeds all baselines in accuracy, including the best full-transmission method AttFusion, while using only 44 KB per frame—a $520{\times}$ reduction from AttFusion’s 22,792 KB. On the real-world DAIR-V2X dataset, JigsawComm achieves the highest mAP@0.5 (0.69), surpassing V2X-ViT (0.68) while requiring only 282 KB ($116{\times}$ less data). Among communication-efficient baselines, JigsawComm outperforms the strongest competitor ERMVP by 3 mAP points on OPV2V and 5 on DAIR-V2X, while using $50{\times}$ and $21{\times}$ less bandwidth, respectively, thanks to the joint encoder–scheduler design eliminating both intra- and inter-agent redundancy simultaneously.
@@
Figure 5: Accuracy vs. required bandwidth. JigsawComm (curve) is flexible via $B$; baselines (points) are fixed.
(a) Accuracy
(b) Feature size
Figure 6: Scalability on OPV2V by #agents. JigsawComm bounds the payload to $\mathcal{O}(1)$.
Accuracy–Bandwidth Tradeoff. Fig. 5 plots JigsawComm’s accuracy as a function of the available bandwidth. On OPV2V, JigsawComm reaches 0.90 mAP@0.5 at only $\sim$}200 Kbps and saturates beyond $\sim$}1,000 Kbps, well within the V2X capacity ($\leq$}20 Mbps)[FCC_DSRC, CV2X_ETSI_EN_303_613]. On DAIR-V2X, the curve saturates at a similarly low bandwidth, confirming that the learned utility proxy successfully identifies the small fraction of features that drive perception accuracy. This property is especially valuable for congested V2X channels, where the scheduler can gracefully degrade by transmitting only the highest-utility cells first. Note that a single trained model serves the entire curve; the budget $B$ acts as a simple “knob” at inference.
Scalability with Number of Agents. Fig. 6 validates $\mathcal{O}(1)$ data-channel scaling on OPV2V. The total feature size grows only slightly from $\approx$}40 KB (N=2N{=}2) to $\approx$}58 KB (N=4N{=}4) due to increased coverage, while accuracy remains high. This sub-linear growth is a direct consequence of the top-1-per-cell policy (Theorem 3.1): as more agents observe the same region, the scheduler selects only the single best contributor, converting redundancy into bandwidth savings. Note that the OPV2V benchmark groups scenes by the number of participating agents rather than incrementally adding agents to a fixed scene. Consequently, differences in scene composition (object density, spatial layout) across groups can cause slight non-monotonicity in the per-group averages (e.g. a small decrease from N=4N{=}4 to N=5N{=}5). The overall sub-linear trend nevertheless holds clearly across the range. We further test on a custom dataset with 10–20 CAVs (Appendix 0.D): the payload saturates beyond $\sim$}15 agents, confirming $\mathcal{O}(1)$ scaling as inter-agent overlap dominates.
@@
End-to-End Efficiency. The average end-to-end latency is $45.82{\pm}26.96$ ms, composed of $28.27{\pm}23.88$ ms computation on a single RTX 4090 and $17.55{\pm}7.14$ ms communication at 20 Mbps, which is well within the 100 ms real-time budget for autonomous driving [ADConstraints]. The meta utility map overhead from all agents is only $0.08{\pm}0.04$ KB per frame, an average of less than $0.03$ KB/agent under FP4.
4.3 Robustness Analysis
Real-world V2X deployments are subject to imperfect localization and communication delays. We evaluate JigsawComm’s robustness to both factors on the OPV2V dataset and compare against ERMVP [Zhang_2024_CVPR] and CoST [tang2025cost]. We further show JigsawComm’s robustness against packet drops in Appendix 0.E.
(a) Localization error
(b) Delay
Figure 7: Robustness on OPV2V. JigsawComm degrades gracefully under localization noise and communication delay.
(a) OPV2V
(b) DAIR-V2X
Figure 8: Top-1 vs. Top-2 agents per cell. Redundant features degrade accuracy, corroborating Theorem 3.1.
Localization Error. We inject Gaussian pose noise with $\sigma\in\{0,0.2,\ldots,1.0\}$ m into each agent’s reported position. As shown in Fig. 7(a), JigsawComm degrades gracefully from 0.90 ($\sigma{=}0$) to 0.74 ($\sigma{=}1.0$ m), a relative drop of 18%, thanks to the RANSAC-based utility map alignment (Sec. 3.6). In contrast, CoST drops from 0.87 to 0.39 (55%). The resilience of JigsawComm stems from the fact that the sparse, high-utility cells used for RANSAC matching are predominantly located on or near objects, providing robust correspondences even under noise. At the practically relevant range of $\sigma{\leq}0.4$ m, JigsawComm’s accuracy remains above 0.85.
Communication Delay. We simulate delays of $\{0,100,\ldots,500\}$ ms by substituting each agent’s current features with those from an earlier timestamp. Fig. 7(b) shows JigsawComm degrades more gracefully: at 200 ms it retains AP@0.5 of 0.72 (vs. 0.59 ERMVP, 0.41 CoST); at 500 ms—equivalent to 5 missed frames at 10 FPS—JigsawComm still achieves 0.63, outperforming both baselines by a large margin. We attribute this resilience to two complementary effects: (i) the scheduler selects only the most informative features, so the fused representation is less sensitive to the staleness of any individual contribution; and (ii) the meta utility map and feature warping via RANSAC compensate for spatial misalignment induced by asynchronous transmission [zhang2023robust].